OpenStudy (anonymous):

given two points M and N on the coordinate plane, find the slope of line MN, and state the slope of the line perpendicular to line MN: M(-2,2) N(4,-4)

OpenStudy (anonymous):

The slope or gradient of the line is: \[((-4)-(2))/((4)-(-2)) = -1\] And the gradient of the perpendicular line is the negative reciprocal, that is: \[-1/-1 = 1\]

OpenStudy (anonymous):

(2--4)/(-2-4)=-1. slope perpendicular is 1.

OpenStudy (anonymous):

the formula to find the slope is below \[ m = \frac{y_{2}-y_{1}} {x_{2}-x_{1}} \] m=slope Just plug your points in and do the math